### v1

# Wormhole Formation with Gravitational Particle Creation Mechanism

Abstract
In this paper, we explore the possibility of the formation of a traversable wormhole geometry in General Relativity sustained by particle creation mechanism. Gravitational particle creation mechanism is a thermodynamic approach to generate repulsive back-reaction gravitational effects through the thermodynamic creation of particles. This repulsive back-reaction pressure can be thought of as a source of sustaining a traversable wormhole. In the first part of this paper, we systematically model a wormhole geometry through the process of particle creation by assuming an inverse square variation law of particle creation pressure within the wormhole geometry, and the shape function for the wormhole is obtained, assuming the redshift function as constant and finite. The obtained shape function is found to adhere to the feasibility conditions (flaring out condition and asymptotic flatness condition) of a viable wormhole geometry. In the second part of the paper, again by assuming a constant and finite redshift function, we studied the radial evolutionary behaviour of the particle creation pressures for three commonly utilized toy shape-functions. The examination of the energy conditions namely the Null Energy Condition (NEC) and the Dominant Energy Condition (DEC) for all these models reveal that the NEC is violated for the shape-function derived from the first principles of particle creation mechanism, as well as for the toy shape functions. The DEC is however found to be satisfied for all the aforementioned cases. In essence, NEC violation occurs through the particle creation phenomena inside the wormhole structure in spite of the existence of non-exotic matter.
Wormholes; Gravitational Particle Creation; Cosmology; Energy Condition
I Introduction
Wormholes at present are interstellar theoretical passages connecting two distant regions in spacetime, which have been intriguing the minds of physicists, since the advancement of general relativity (GR). These interesting structures are one of the many possible solutions to the Einstein field equations (EFE), which presents itself as a fascinating topic for systematic investigations in different technical frameworks.
A class of physically reasonable traversable wormhole structures are first described by Morris and Thorne in GR to explore the possibility of interstellar space travel without violating the laws of special relativity (SR)[1]. Earlier studies demonstrated the possibility of traversable wormholes in GR [2, 3]. A more technical definition of a traversable wormhole is a topological rupture in the fabric of spacetime that allows for traversable interstellar connections as well as other temporal links between distant spacetime points. A traversable wormhole with exotic matter leads one to expect that the energy conditions are violated in the wormhole. For instance, it was argued that Casimir energy can be a possible source to sustain a traversable wormhole geometry [4, 5]. Several works validating the requirement of violation of the energy conditions for the existence of wormholes can be found in [6, 7, 8, 9, 10, 11, 12] .
Observational data suggests that a significant portion of the total matter content is probably occupied by dark energy, which is responsible for the accelerated expansion of the Universe. As a basis of this observational result, the background dynamics are provided by the GR (as well as modified theories of gravity). Detailed information can be found in references [13, 14]. Generically, a cosmological fluid capable of generating negative pressure can be accounted as a possible explanation for the so-called dark energy. In the literature, there are several models of dark energy, which theoretically explain the accelerated expansion of the Universe with relevance to observed cosmological data [15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. In this context, a possible explanation for the observed accelerated expansion of the Universe, gravitational particle creation, is a theory based on non-equilibrium thermodynamics proposed by Prigogine [25], provides a mechanism capable of generating negative pressure in the Universe, through the production of particles involving thermodynamic processes at the expense of the gravitational field. Its primary advantage is that the negative pressure as a result of the creation of particles (also called back-reaction) is naturally accommodated by the second law of thermodynamics. Since the particle creation process is macroscopically described by a negative back-reaction pressure of quantum origin, these models have been studied as a possible means to explain the observed phenomena of the dark sector of cosmology [26, 27, 28]. Motivated by the aforementioned context, we suggest that the negative pressure produced as a result of the particle creation processes when incorporated into the energy-momentum tensor of the EFE, can be expected to contribute to the violation of the energy conditions required for the formation of wormholes. Therefore, we plan the paper in the following way: In section II, we develop the theoretical setting of the wormhole in the presence of gravitational particle creation and hence derive the shape function that describes the wormhole geometry. In section III, we assume different toy models of some viable wormhole geometry and obtain how particle creation pressure evolves inside the wormhole geometry. In section IV, we derive the embedding function and provide a 3D visualization of the wormhole shape function obtained through the requirements of particle creation. In section V, we finally discuss and conclude the results of our study.
II Wormhole Solution from Particle Creation Mechanism
Let us assume the spherically symmetric wormhole spacetime given by the metric
$ds^{2}=-e^{2\Phi(r)}dt^{2}+\frac{dr^{2}}{1-b(r)/r}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right).$ (1)
The metric functions $\Phi(r)$ and $b(r)$ are general arbitrary functions of the radial coordinate $r$ within the context of a wormhole spacetime. The redshift function, denoted as $\Phi(r)$, characterizes the gravitational redshift effect, while the shape function, denoted as $b(r)$, provides information about the geometry of the wormhole. It is essential to note that the radial coordinate $r$ exhibits non-monotonic behavior, wherein it decreases from infinity to a minimum value $r_{0}$, representing the location of the throat of the wormhole. At the throat, $b(r_{0})=r_{0}$. Subsequently, the radial coordinate increases from $r_{0}$ to infinity.
[…]
In this paper, we address wormhole formation within the context of irreversible matter creation formalism and thermodynamics of open systems in cosmology, as described by Prigogine [25]. This formalism views the Universe as an open system, with particle formation described by an adaptation of the energy-momentum tensor that incorporates a matter creation term into conservation laws. Consider an open system of volume $V$, containing $N(t)$ particles with energy density $\rho$ and thermodynamic pressure $p$. The thermodynamic conservation equation in this system is
$d(\rho V)=dQ-pdV+\frac{h}{n}d(nV),$ (2)
In this equation, $dQ$ represents the heat absorbed by the system during time $dt$, $h=\rho+p$ is the enthalpy per unit volume, and $n=N/V$ is the particle density. In open systems, thermodynamic energy conservation includes the production and annihilation of matter, unlike isolated or closed systems where the number of particles is constant.
At this point, we use the formalism of irreversible matter creation from the thermodynamics of open systems to cosmology by considering a homogeneous and isotropic Universe as an open system with volume $V$, $N(t)$ particles, energy density $\rho$, and thermodynamic pressure $p$, as represented by the FLRW metric
$ds^{2}:=-dt^{2}+a^{2}(t)[dr^{2}+r^{2}d\Omega^{2}],\quad d\Omega^{2}=d\theta^{2}+sin^{2}\theta d\phi^{2}.$ (3)
In a homogeneous and isotropic Universe, the volume of the Universe may be represented using the scale factor $a$ as $V=a^{3}(t)$. Equation (2) may be recast in terms of total time derivatives of physical quantities as follows [25]:
$\frac{d}{dt^{2}}(\rho a^{3})+p\frac{d}{dt}a^{3}=\frac{dQ}{dt}+\frac{\rho+p}{n}\frac{d}{dt}(na^{3}).$ (4)
. As homogenous systems do not receive heat, assuming the Universe is homogeneous, we can speculate that the heat received remains constant over time ($dQ/dt=0$). This result enables us to reformulate equation (4) in an equivalent form:
$\dot{\rho}+3H(\rho+p)=\frac{\rho+p}{n}(\dot{n}+3Hn).$ (5)
In an open thermodynamical system, the non-conservation of the total rate of change of the number of particles, $N=na^{3}$, with comoving volume $a^{3}$ and $n$ being the number density of particle production, yields [29, 30]
$\dot{n}+3Hn=n\Gamma,$ (6)
where $\Gamma$ is a particle creation rate that has negative and positive phases. Negative $\Gamma$ represents the particle destruction and positive $\Gamma$ describes the annihilation of particles.
The equation relating the creation rate $\Gamma$ to the creation pressure $\Pi$ is
$\Pi:=-\frac{\Gamma}{3H}(\rho+p).$ (7)
In this work, we intend to investigate the existence of a traversable wormhole within a framework where the negative pressure required for sustaining the wormhole is provided by the particle creation mechanism given by [29, 30, 31, 32, 33, 34]
$\Pi:=-\frac{\Gamma}{3H_{0}}(\rho+p)=-\frac{\Gamma\rho}{3H_{0}}(1+\omega),$ (8)
where $\Gamma$ is the source function, $\beta$ is an arbitrary constant in the range $[0,1]$ and $\omega$ is the barotopic equation of state (EoS) parameter. Here, as we are investigating the wormhole solution in the present era, we have set $H=H_{0}$, which is the present value of the Hubble parameter. Moreover, we assume that the matter inside the wormhole is perfect fluid. In this work, we consider specifically that the perfect fluid follows a dust-like equation of state, i.e $\omega=0$.
In this work, we shall work with the commonly used form of the particle creation rate given by [31]
$\Gamma=3\beta H_{0},$ (9)
The set of field equations necessary for our study then takes the form
$\rho(r)$ $=\frac{b^{\prime}}{8\pi r^{2}},$ (10)
$p_{r}(r)$ $=-\frac{b}{r^{3}}-\Pi(r),$
$p_{t}(r)$ $=-\left(1-\frac{b}{r}\right)\left(\frac{b^{\prime}r-b}{2r^{2}(r-b)}\right)-\Pi(r).$
where $\rho(r)$ is the energy density, $p_{r}(r)$ and $p_{t}(r)$ are the radial and tangential pressures respectively which are mutually orthogonal. $\Pi(r)$ is the particle creation pressure introduced into the energy-momentum tensor $T^{\mu}_{\nu}=\text{diag}[-\rho(r),p_{r}(r)+\Pi(r),p_{t}(r)+\Pi(r),p_{t}(r)+\Pi(r)]$. Subsituting $\rho$ and (9) into equation (8) we obtain
$\Pi(r)=-\beta(1+\omega)\frac{b^{\prime}(r)}{8\pi r^{2}},$ (11)
To obtain the explicit form of the shape function $b(r)$, the particle creation pressure should be specified in terms of the radial coordinate $r$. There may be different functional variations to how $\Pi(r)$ may evolve with $r$, but from a phenomenological standpoint, let us propose an inverse-square variation law of the particle creation pressure that varies with the radial coordinate $r$ as
$\Pi=-\frac{K}{r^{2}},$ (12)
where $K$ is a proportionality constant, that accounts for the dimensions of $\Pi(r)$. This form of variation of particle creation pressure implies that the particle creation pressure is maximum at the throat ($r=r_{0}$) and decreases inverse-squarely with both radial directions. Nevertheless, one may still choose a general power-law variation ($\propto-1/r^{\alpha}$) as well. We considered the specific case by setting $\alpha=2$ for mathematical simplicity along to keep the model limited to the least number of free parameters.
Now, substituting equation (12) in equation (11), we obtain the shape function for our model as
$b(r)=r_{0}+\frac{8\pi K}{\beta(1+\omega)}(r-r_{0}),$ (13)
The necessary conditions for a viable wormhole geometry are
  1. 1.
    Throat condition: At the throat $r_{0}$, $b\left(r_{0}\right)=r_{0}$ and $01-\frac{b(r)}{r}>$ 0 outside the throat.
  2. 2.
    Flaring out condition: $b^{\prime}\left(r_{0}\right)<1$ i.e.
    $\frac{b(r)-rb^{\prime}(r)}{b^{2}(r)}>0$
    , where ’ represents derivative w.r.t. $r$.
  3. 3.
    Asymptotically Flatness condition:
    $\frac{b(r)}{r}\rightarrow 0$
    as $r\rightarrow\infty.$
Fig. 1 shows that the flareout and asymptotic conditions are fulfilled, with three sets of the parameter values $\beta$ and $K$. We can see clearly that the shape function (13) describes a viable wormhole geometry.
Refer to captionRefer to caption
Figure 1: The flareout and asymptotic conditions are shown for a few combinations of the model parameter $\beta$ and the proportionality constant $K$. Here throat radius $r_{0}$ is assumed to be 1.
The shape function given by equation (13) determines the configuration of the wormhole which is sustained by the negative particle creation pressure. Interestingly, the wormhole does not need an ad-hoc introduction of exotic fluids to account for the negative pressure required to sustain the wormhole geometry. To explore the physical nature of the matter content inside the wormhole geometry, we shall make use of the energy conditions. A generic aspect of wormhole systems is the violation of the null energy condition (NEC) (i.e. $\rho+p_{r}\geq 0$ or $\rho+p_{t}(r)\geq 0$ in terms of radial and tangential pressures respectively) at the throat of the wormhole structure. On the other hand, for some specific models keeping valid the NEC while keeping the dominant energy condition (DEC) ($\rho-p_{r}(r)\geq 0$ or $\rho-p_{t}(r)\geq 0$) violated at the throat is also possible [35, 36, 37, 38, 39]. In this work, we shall construct traversable wormhole structures with the violation of the NEC as well as DEC at the throat of the wormhole, simultaneously keeping the energy condition $\rho\geq 0$. This turns out, as a result of the violation of the weak energy condition (WEC) ($\rho\geq 0,\rho+p_{r}(r)\geq 0\text{ or }\rho+p_{t}(r)\geq 0$). From equation (10), given the shape function (13), one may obtain the energy density, radial and tangential pressures given by
$\rho(r)=\frac{K}{\beta r^{2}(\omega+1)},$ (14)
$p_{r}(r)=\frac{K\left(r-\frac{8\pi(r-r_{0})}{\beta(\omega+1)}\right)-r_{0}}{r^{3}},$ (15)
$p_{t}(r)=\frac{2Kr-\frac{8\pi Kr_{0}}{\beta\omega+\beta}+r_{0}}{2r^{3}},$ (16)
To study the energy conditions for the wormhole geometry (13), let us calculate the quantities $\rho+p_{r}(r)$, $\rho+p_{t}(r)$, $\rho-p_{r}(r)$ and $\rho-p_{t}(r)$ given by
$\rho+p_{r}(r):=\frac{Kr(\beta\omega+\beta-8\pi+1)+8\pi Kr_{0}-\beta r_{0}(\omega+1)}{\beta r^{3}(\omega+1)},$ (17)
$\rho+p_{t}(r):=\frac{2K\left(\frac{r-4\pi r_{0}}{\beta\omega+\beta}+r\right)+r_{0}}{2r^{3}},$ (18)
and
$\rho-p_{r}(r):=\frac{\frac{K(8\pi r+r-8\pi r_{0})}{\beta(\omega+1)}-Kr+r_{0}}{r^{3}}.$ (19)
$\rho-p_{t}(r):=-\frac{2K\left(r-\frac{r+4\pi r_{0}}{\beta\omega+\beta}\right)+r_{0}}{2r^{3}}.$ (20)
Figure 2: The NEC and DEC for the (13) is shown for three values of $K$ and $\beta$
Fig. 2 displays the behaviour of the NEC and DEC for different combinations of the model parameters. The NEC in terms of radial pressure shows clear violation at the throat but obeys at far away regions from the throat, whereas the DEC is obeyed at the throat as well as at far away regions. However, in terms of the tangential pressure, the NEC is obeyed at the throat as well as in asymptotic regions but DEC is violated at the throat and satisfied at the asymptotic regions. Also as $\rho\geq 0$, one can say WEC is violated for the radial pressure case and obeyed for the tangential pressure one.
III Toy Shape Functions and Evolution of Particle Creation Pressure
In the previous section, we utilized a phenomenological variation law of the particle creation pressure and found a shape function corresponding to that scenario. In this section, let us assume some specific shape functions and study the evolution of the particle creation pressure.
III.1 Shape function 1: $b(r)=\frac{r_{0}\tanh(\epsilon r)}{\tanh(\epsilon r_{0})}$
Let us now assume a hyperbolic shape function given by
$b(r)=\frac{r_{0}\tanh(\epsilon r)}{\tanh(\epsilon r_{0})},$ (21)
where $\epsilon$ is a constant parameter. This form of shape function has been studied in different contexts like Lyra manifold [40] and $f(R)$ gravity [41].
The energy density with this shape function is
$\rho=\frac{r_{0}\epsilon\operatorname{sech}^{2}(r\epsilon)\coth(r_{0}\epsilon)}{8\pi r^{2}},$ (22)
which clearly is a positive function for all values of $r$. From Fig.3, it is seen that the shape function obeys the viability conditions mentioned in section II for different values of $\epsilon$.
Refer to caption Refer to caption
Figure 3: The flaring out and asymptotic flatness condition for the shape function (21) is shown for different combinations of the model parameters.
The radial pressure $p_{r}$ and tangential pressure $p_{t}$ in this case take the form
$p_{r}=\frac{r_{0}\operatorname{sech}^{2}(r\epsilon)\coth(r_{0}\epsilon)(\beta r(\omega+1)\epsilon-4\pi\sinh(2r\epsilon))}{8\pi r^{3}},$ (23)
$p_{t}=\frac{r_{0}\operatorname{sech}^{2}(r\epsilon)\coth(r_{0}\epsilon)(r\epsilon(\beta\omega+\beta-4\pi)+2\pi\sinh(2r\epsilon))}{8\pi r^{3}}.$ (24)
The particle creation pressure for the shape function (21) can be obtained from equation (11)
$\Pi(r)=-\frac{\beta r_{0}(\omega+1)\epsilon\operatorname{sech}^{2}(r\epsilon)\coth(r_{0}\epsilon)}{8\pi r^{2}}$ (25)
Also, the quantities related to the energy conditions are found to be
$\rho+p_{r}:=\frac{r_{0}\operatorname{sech}^{2}(r\epsilon)\coth(r_{0}\epsilon)(r\epsilon(\beta\omega+\beta+1)-4\pi\sinh(2r\epsilon))}{8\pi r^{3}},$ (26)
$\rho+p_{t}:=\frac{r_{0}\operatorname{sech}^{2}(r\epsilon)\coth(r_{0}\epsilon)(r\epsilon(\beta\omega+\beta-4\pi+1)+2\pi\sinh(2r\epsilon))}{8\pi r^{3}},$ (27)
$\rho-p_{r}:=\frac{r_{0}\operatorname{sech}^{2}(r\epsilon)\coth(r_{0}\epsilon)(4\pi\sinh(2r\epsilon)-r\epsilon(\beta\omega+\beta-1))}{8\pi r^{3}},$ (28)
$\rho-p_{t}:=-\frac{r_{0}\operatorname{sech}^{2}(r\epsilon)\coth(r_{0}\epsilon)(r\epsilon(\beta\omega+\beta-4\pi-1)+2\pi\sinh(2r\epsilon))}{8\pi r^{3}},$ (29)
Figure 4: NEC and DEC for the shape function (21) for different combination of the model parameters
From Fig. 4, it is observed that the NEC is violated in terms of radial pressure shown at the throat but tends to obey at the asymptotic limit, whereas the DEC is obeyed at the throat as well as at the asymptotic limit. In terms of the tangential pressure, the NEC is obeyed at the throat as well as at asymptotic regions but DEC is obeyed at the throat for the parameter values $(\beta,\epsilon)=(0.2,0.10)\text{ and }(0.5,0.15)$, however, crosses into the violation region at around $r=3$ and $r=1.6$ respectively for the ranges we considered. For $(\beta,\epsilon)=(0.7,0.20)$, the DEC is always violated. Also, as $\rho\geq 0$, WEC is violated for the radial pressure case and obeyed for the tangential pressure one. Moreover, from equation (25) and Fig. 5, it is evident that the particle creation pressure is negative at the throat all the way to $r\to\infty$, which is required and expected for backreaction pressure responsible for sustaining a wormhole geometry.
Figure 5: Particle creation pressures with shape function (21)
III.2 Shape function 2: $b(r)=r_{0}^{n}r^{1-n}$
Another viable wormhole is a power-law type shape function given by
$b(r)=r_{0}^{n}r^{1-n},$ (30)
where $n$ is a model parameter and $r_{0}$ is the throat radius. This type of shape function has been studied in [41], in the context of $f(R)$ gravity reconstruction for given choices of wormhole shape functions. Fig. 6 clearly shows the required criteria satisfied for viability for different combinations of the associated model parameters.
Refer to captionRefer to caption
Figure 6: The flaring out and asymptotic flatness condition for the shape function (30) is shown for different combinations of the model parameters.
The energy density for this particular shape function takes the form
$\rho=\frac{(1-n)r^{-n-2}r_{0}^{n}}{8\pi},$ (31)
The radial and tangential pressure can be calculated as
$p_{r}(r)=\frac{\beta(1-n)(\omega+1)r^{-n-2}r_{0}^{n}}{8\pi}-r^{-n-2}r_{0}^{n},$ (32)
$p_{t}(r)=\frac{r^{-n-2}r_{0}^{n}(\beta(\omega+1)-\beta n(\omega+1)+4\pi n)}{8\pi},$ (33)
The particle creation pressure can be obtained as
$\Pi(r)=-\frac{\beta(1-n)(\omega+1)r^{-n-2}r_{0}^{n}}{8\pi}.$ (34)
Also, we have the quantities related to the energy conditions
$\rho+p_{r}:=\frac{r^{-n-2}r_{0}^{n}(\beta\omega+\beta-n(\beta\omega+\beta+1)-8\pi+1)}{8\pi},$ (35)
$\rho+p_{t}:=\frac{r^{-n-2}r_{0}^{n}(\beta\omega+\beta+n(-\beta(\omega+1)+4\pi-1)+1)}{8\pi},$ (36)
$\rho-p_{r}:=\frac{r^{-n-2}r_{0}^{n}(\beta(-\omega)-\beta+n(\beta\omega+\beta-1)+8\pi+1)}{8\pi},$ (37)
$\rho-p_{t}:=-\frac{r^{-n-2}r_{0}^{n}(\beta\omega+\beta+n(\beta(-\omega)-\beta+4\pi+1)-1)}{8\pi},$ (38)
Figure 7: NEC and DEC for the shape function (30)
Refer to caption     Refer to caption
Figure 8: Particle creation pressures with shape functions (30) (left) and (39) (right)
Fig. 7, shows that the NEC is violated in terms of radial pressure at the throat but tends to obey as $r\to\infty$, whereas the DEC is obeyed at the throat as well as at $r\to\infty$. In terms of the tangential pressure, the NEC is obeyed at the throat as well as at asymptotic regions but DEC is violated at the throat and obeyed at $r\to\infty$. Also, WEC is violated in terms of the radial pressure but obeyed in terms of the tangential pressure. Moreover, from equation (34) and Fig. 8 (left), it is seen that the particle creation pressure is negative at the throat to $r\to\infty$, but with an exception for the shape function (39) in that the particle creation pressure does not tend to 00 as $r\to\infty$ like all the other cases, but tends to higher negative values at regions far away from the throat.
III.3 Shape function 3: $b(r)=r\exp[\Omega(r-r_{0})]$
Finally, we consider an exponential-type shape function given by
$b(r)=r\exp[\Omega(r-r_{0})],$ (39)
where $\Omega$ is the only model parameter and $r_{0}$ being the throat radius. This type of shape function has been considered widely in Ref. [42, 43, 44] in both GR and modified gravity. Fig. 9 shows the validity of the flaring-out and asymptotic flatness for different combinations of the associated model parameters.
Refer to captionRefer to caption
Figure 9: The flaring out and asymptotic flatness condition for the shape function (39) is shown for different combinations of the model parameters.
The energy density corresponding to this shape function is
$\rho=\frac{e^{-(\Omega(r-r_{0}))}-r\Omega e^{-(\Omega(r-r_{0}))}}{8\pi r^{2}},$ (40)
The radial and tangential pressures are obtained as
$p_{r}(r)=\frac{\beta(\omega+1)\left(e^{-(\Omega(r-r_{0}))}-r\Omega e^{-(\Omega(r-r_{0}))}\right)}{8\pi r^{2}}-\frac{e^{-(\Omega(r-r_{0}))}}{r^{2}},$ (41)
$p_{t}(r)=\frac{e^{\Omega(r_{0}-r)}(4\pi r\Omega-\beta(\omega+1)(r\Omega-1))}{8\pi r^{2}},$ (42)
In this model, the particle creation pressure can be expressed as
$\Pi(r)=-\frac{\beta(\omega+1)\left(e^{-(\Omega(r-r_{0}))}-r\Omega e^{-(\Omega(r-r_{0}))}\right)}{8\pi r^{2}}.$ (43)
The quantities associated with the energy conditions are calculated as
$\rho+p_{r}(r):=-\frac{e^{\Omega(r_{0}-r)}((\beta\omega+\beta+1)(r\Omega-1)+8\pi)}{8\pi r^{2}},$ (44)
$\rho+p_{t}:=\frac{e^{\Omega(r_{0}-r)}(-\beta(\omega+1)(r\Omega-1)+(4\pi-1)r\Omega+1)}{8\pi r^{2}},$ (45)
$\rho-p_{r}:=-\frac{e^{\Omega(r-r_{0})}((\beta\omega+\beta-1)(r\Omega+1)+8\pi)}{8\pi r^{2}},$ (46)
$\rho-p_{t}:=\frac{e^{\Omega(r_{0}-r)}((\beta\omega+\beta-1)(r\Omega-1)+8\pi)}{8\pi r^{2}},$ (47)
Figure 10: NEC and DEC for the shape function (39) for different combinations of model parameters.
From Fig. 10, we see NEC violation in terms of radial pressure at the throat but is obeyed at $r\to\infty$, whereas the DEC is obeyed at the throat as well as at $r\to\infty$. In terms of the tangential pressure, the NEC is satisfied at the throat as well as at asymptotic regions but DEC remains violated at the throat but obeyed at $r\to\infty$. Moreover, WEC is violated for the radial pressure case and satisfied for the tangential pressure one. Also, from equation (43) and Fig. 8 (right), it is seen that the particle creation pressure is negative at the throat and tends to zero at $r\to\infty$.
IV 3D Embedding
To realize the embedding function of a wormhole, one may choose the equatorial slice $\theta=\pi/2$ and $t=$constant. This reduces the metric (1) to the form
$ds^{2}=\left(1-\frac{b(r)}{r}\right)^{-1}dr^{2}+r^{2}d\phi^{2}.$ (48)
The metric (48) is equivalent to
$ds^{2}=dz^{2}+dr^{2}+r^{2}d\phi^{2},$ (49)
in cylindrical coordinates. The embedded surface may be related to a 3D-Euclidean surface defined by $z=z(r)$, the metric (49) can be further expressed in the following form
$ds^{2}=\left[1+\left(\frac{dz}{dr}\right)^{2}\right]dr^{2}+r^{2}d\phi^{2},$ (50)
which can now be compared with equation (48) to obtain
$\frac{dz}{dr}=\pm\left(\frac{r}{b(r)}-1\right)^{-1/2},$ (51)
Equation (51) can be solved to obtain a functional form of $z(r)$. The shape function obtained from the physical requirements of particle creation given by (13) is shown in the Fig. 11 for $\beta=0.7$ and $K=0.016$.
Figure 11: The 3D embedding of the wormholes are shown for the shape functions (13) the throat is assumed at $r_{0}=1$ and the values of $\beta$ and $K$ is taken to be $0.7$ and $0.016$ respectively.
V Conclusion
The concept of gravitational particle creation phenomena of open thermodynamic cosmological systems, was proposed as a potential explanation to the late-time accelerated expansion of the Universe and inflationary expansion during the early stages of the Universe. This mechanism involves the generation of repulsive gravitational effects, resulting in negative pressure, which could explain the accelerated expansion of the Universe in its later stages. Gravitational particle creation allows for speculation about violations of the NEC, which is crucial for the maintenance of a wormhole structure. This study explores the hypothesis that the geometry of a wormhole can be supported by the negative pressure created by particles. The initial section of the paper focuses on deriving the shape function of a wormhole based on a specific form of particle creation pressure. Specifically, we proposed an inverse-squared variation of particle creation pressure with the radial coordinate, with a source function that is linear in the Hubble parameter at the current time, denoted by $\Gamma=3\beta H_{0}$ with $\beta$ as a free parameter. Through out the paper, owing to the finiteness of the redshift function, we considered a constant redshift function for the analysis. The shape function obtained seems to adhere to the required criteria for a viable wormhole structure and violates the NEC with a positive energy density. In the later part of the paper, in the context of particle creation mechanism we obtained the radial variation of energy density, radial and tangential pressure, and the particle creation pressure within the wormhole geometry, by assuming a few well known shape functions. The radial NEC is found to be violated at the throat in all the cases, which resonates with the behaviour of so-called exotic matter but without the need of introducing explicit form of exotic matter density like chaplygin gas [45, 39, 46, 47, 48, 49, 50, 51] or phantom matter [52, 53, 54, 55, 56, 57, 58, 59, 60]. Finally, the 3D embedding for the wormhole solution obtained from the particle creation phenomena is also carried out.
This study explores the possibility for developing a viable wormhole geometry supported by the back reaction pressure generated from the irreversible particle creation, within an open thermodynamic system. Nevertheless, our work does not address the effect of particle creation on particle geodesics within the wormhole structure or related aspects, as it is beyond the current study’s scope. This prospect will be further investigated in the future work.